A company enters into long futures contracts to buy 50,000 bushels of corn in May for 411 cents per bushel. The initial margin rate is $1,375 per contract and the maintenance margin is $1,250 per contract (one contract is 5,000 bushels). What price would lead to a margin call? O A 436 cents or above O B 413.50 cents or above O C 408.50 cents or below O D 386 cents or belowIn the previous situation (question - 2.6), under what circumstances up to $3,000 can be withdrawn from the margin account? O A If price goes below 405 cents O B If price stays 411 cents O C If price goes above 417 cents O D If price goes above 471 centsPlease show how to get these answers.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut ta copy ▼ E AutoSum Calibri ー E ゴWrap Text General в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard BK140 BD Font Alignment Number Styles Cells Edting BE BF BG BH BI BJ BK BL BM BN BO 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 ー4 ‘ ト 125 MARGIN CALL WILL BE THERE, WHEN THE VALUE OF CONTRACT FALLS BY 1375-1250 = AS CONTRACT IS OF 5000 BUSH ELS, THE PRICE PER BUSHEL SHOULD FALL BY = 0.025 IN DOLLARS 125/5000 2.5 IN CENTS SO PRICE AT MARGIN CALL-411-2.5- 408.5 NS1: C: INITIAL MARGIN = 1375 1250 MAINTENANCE MARGIN = IF WE WANT TO WITHDRAW 3000 AND AS THERE ARE 10 CONTRACTS, PER CONTRACT, THE WITHDRAWAL IS = 300 SO MARGIN SHOULD INCREASE BY- 300 AS CONTRACT IS OF 5000 BUSHELS, THE PRICE PER BUSHEL SHOULD INCREASE BY = 0.06 IN DOLLARS 300/5000 6 IN CENTS SO PRICE WILL BE = 411 + 6 = 417 ANS 2: C OCF CASH FLOW STAT -FCF -ARM MSI confidence!NDEX CURRENCY LOAN CROSS OVER HEDGING MONEY MARKET -GM HPR AM TAX GROWTH SUST-PREF AND DIV POLICY 08:41 31-01-2019